### v1

@@ 3 Exact operator inference
In this work, we propose a novel method to generate snapshots that is a priori guaranteed to result in a full rank matrix ${\mathbf{P}}$. Our method has two key ingredients. First, in Section 3.1, we leverage the knowledge about how the matrix ${\mathbf{P}}$ depends on the ROM states $\breve{\bm{x}}_{k}$ and input signals ${\bm{u}}_{k}$ via (12) and (7) to find pairs of ROM states $\bar{\bm{x}}_{s}\in\mathbb{R}^{n}$ and input signals $\bar{\bm{u}}_{s}\in\mathbb{R}^{N_{u}}$, $s=1,\dots,{n_{f}}$ that ensure full rank of ${\mathbf{P}}$. Second, in Section 3.2, we generate time derivative snapshots $\dot{\bar{\bm{x}}}_{s}$ cooresponding to these rank-ensuring pairs $(\bar{\bm{x}}_{s},\bar{\bm{u}}_{s})$ by simulating ${n_{f}}$ trajectories for only a single time step each, with initial conditions $\bar{\bm{x}}_{s}$ and input signals $\bar{\bm{u}}_{s}$.
@@ 4.2 Ice sheet model
For ROM dimension $n=7$, the proposed method performs ${n_{f}}=3087$ FOM time steps. This number is larger than the $2001$ FOM time steps with re-projection reported in [2]. However, the authors also report that this budget of snapshots is depleted before the inference of the ROM operators of dimension $n=6$ is completed. Hence, only ROM operators of dimension up to $n=5$ are computed with re-projected snapshots and exact operator inference. The ROM operators of dimensions $n=6$ and $n=7$ are partially computed with regularized operator inference and snapshot data polluted with non-Markovian terms. Consequently, the inferred operators of dimension $n=6$ and $n=7$ are not reconstructed exactly. The operators of dimension $n\leq 5$ are indeed exactly reconstructed, but the used $2001$ snapshots are more than the ${n_{f}}=365$ snapshots used by our method for dimension $n=5$.
@@ Appendix A Proof of full rank of matrix ${\mathbf{P}}$
For any $n\in\mathbb{N}^{+}$, $l\in\mathbb{N}_{0}$, for any $\mathcal{I}\subset\mathbb{N}_{l}$, there exists a unique polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}}$ such that
@@
The special case of degree sets without gaps, (i.e. $\mathcal{I}=\mathbb{N}_{l}$), is covered by an existing result on existence and uniqueness of an interpolating multivariate polynomial [17] described in Section A.1. In Section A.2, we use this result to prove Theorem A for the special case that $\mathcal{I}$ contains only one element. Hence, we can find for all $i\in\mathcal{I}$ a homogeneous polynomial $p_{i}\in\mathcal{P}^{n}_{\{i\}}$ that interpolates in the nodes in $\mathcal{X}^{i}$. To extend this result to the general case where $\mathcal{I}$ contains multiple elements, we combine such homogeneous polynomials via the following algorithm.
Algorithm 2 Gappy multivariate polynomial interpolation
1:procedure getInterpolant(dimension $n$, degree set $\mathcal{I}$, interpolation values $b_{j}$)
2:     $p\leftarrow 0$
3:     for ${i}\in\mathcal{I}$ ordered in descending order do
4:         Find $p_{i}\in\mathcal{P}^{n}_{\{{i}\}}$ that satisfies $p_{i}({\bm{q}}_{j})=b_{j}-p({\bm{q}}_{j})$ for all ${\bm{q}}_{j}\in\mathcal{X}^{i}$
5:         Find $\tilde{p}_{i}$ such that $\tilde{p}_{i}({\bm{q}}_{j})=1$ for all ${\bm{q}}_{j}\in\mathcal{X}^{i}$ and $\tilde{p}_{i}({\bm{q}}_{j})=0$ for all ${\bm{q}}_{j}\in\mathcal{X}^{s}$ for all $s\in\mathcal{I}$ with $s>i$
6:         $p\leftarrow p+\tilde{p}_{i}p_{i}$
7:     end for
return
$p$
8:end procedure
After each iteration $i$ of the for-loop in line 3, the polynomial $p$ satisfies $p({\bm{q}}_{j})=b_{j}$ for all $s\in\mathcal{I}$ with $s\geq i$. Hence, after completion of the for-loop, $p$ satisfies the interpolation problem (39). In addition, we can ensure that $p\in\mathcal{P}^{n}_{\mathcal{I}}$ by restricting the spaces in which the polynomials $\tilde{p}_{i}$ live. The details of this restriction are described in Section A.3. In Section A.4, we combine all results to prove Theorem A.
A.1 Multivariate polynomial interpolation without gaps
@@
A.2 Homogeneous multivariate polynomial interpolation
@@
for arbitrary interpolation values $\hat{f}_{i}$ and the interpolation nodes $\hat{\bm{q}}_{i}$ in $B(l,n-1)$.
@@
A.3 Specific gappy multivariate polynomial interpolation
To ensure that the polynomial $p$ returned by Algorithm 2 is an element of $\mathcal{P}^{n}_{\mathcal{I}}$, we choose $\tilde{p}_{i}$ such that $\tilde{p}_{i}p_{i}\in\mathcal{P}^{n}_{\mathcal{I}}$ for all $i\in\mathcal{I}$. Concretely, we choose $\tilde{p}_{i}\in\mathcal{P}^{n}_{\mathcal{I}^{-i}}$ with $\mathcal{I}^{-i}:=\{s-i|s\in\mathcal{I}\text{ with }s>i\}$ for all $i\in\mathcal{I}$. The following lemma states that there exist polynomials $\tilde{p}_{i}$ in these spaces that satisfy the specific interpolation problems described in line 5 of the algorithm.
Lemma A.1 (Specific gappy multivariate polynomial interpolation).
For any $n\in\mathbb{N}^{+}$, $l\in\mathbb{N}_{0}$, for any $\mathcal{I}\subset\mathbb{N}_{l}$ and for any $i\in\mathcal{I}$, there exists a polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}^{-i}}$ such that
(54) $p({\bm{q}}_{j})$ $=0,\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{s}\text{ for all }s\in\mathcal{I}\text{ with }s>i,\text{ and }$
(55) $p({\bm{q}}_{j})$ $=1,\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i}.$
To prove this lemma, we make the Ansatz that the polynomial $p$ equals a univariate polynomial $\hat{p}$ of the sum of the entries $q_{k}$ of its input ${\bm{q}}$, so $p({\bm{q}})=\hat{p}(\sum_{k=1}^{n}q_{k})$ for all ${\bm{q}}\in\mathbb{R}^{n}$. Since all ${\bm{q}}\in\mathcal{X}^{i}$ satisfy $\sum_{j=1}^{n}q_{j}=i$ for all $i\in\mathbb{N}_{0}$, it remains to prove the univariate case stated in the following lemma.
@@ A.3 Specific gappy multivariate polynomial interpolation
For any $l\in\mathbb{N}_{0}$, for any $\mathcal{I}\subset\mathbb{N}_{l}$ and for any $i\in\mathcal{I}$, there exists a polynomial $p\in\mathcal{P}^{1}_{\mathcal{I}^{-i}}$ such that
(56) $p(s)$ $=0,\hskip 10.00002pt\text{ for all }s\in\mathcal{I}\text{ with }s>i,\text{ and }$
(57) $p(i)$ $=1.$
To prove this lemma, we show in Lemma A.3 that for fixed $l\in\mathbb{N}_{0}$, the existence of a polynomial that solves the specific interpolation problem with interpolation values 0 and 1 as described above implies the existence of a polynomial that solves the general interpolation problem with arbitrary interpolation values. We use this result to prove Lemma A.2 with a case discrimination: In Lemma A.5, we consider the case $i\neq 0$ and in Lemma A.7 the case $i=0$.
@@
for any interpolation values $f_{j}\in\mathbb{R}$.
Proof A.4.
@@
Proof A.6.
The proof proceeds by induction over $l$.
  • Induction start: For $l=0$, $p(x)=1$ satisfies the interpolation properties for the only existing set =$\mathcal{I}=\{0\}$.
  • Induction statement: For fixed $l$, there exists $p\in\mathcal{P}^{1}_{\mathcal{I}}$ such that the interpolation properties (58) and (59) hold for any $\mathcal{I}\in\mathbb{N}_{l}$ and any pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}$.
  • Induction step: We consider a set $\mathcal{I}\subset\mathbb{N}_{l+1}:=\{0,1,\dots,l,l+1\}$ and the pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}^{+}$. If $l+1\notin\mathcal{I}$, then the induction statement directly applies. Otherwise, if $l+1\in\mathcal{I}$, then the induction statement holds for $\mathcal{I}^{-}:=\mathcal{I}\setminus l+1$ and the nodes $x_{1},\dots,x_{{n_{\mathcal{I}}}-1}$. Hence, Lemma A.3 implies that there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
    for any interpolation values $f_{i}\in\mathbb{R}$. Consequently, there exists $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
    satisfies (58). By Descartes’ rule of signs, $\hat{p}$ has at most ${n_{\mathcal{I}}}-1$ positive roots, hence $\hat{p}(x_{n_{\mathcal{I}}})\neq 0$. Consequently, we can normalize $\hat{p}$ to find the desired polynomial,
@@
(70) $\bar{p}(x_{j})=-1\hskip 10.00002pt\text{ for }j=1,\dots,{n_{\mathcal{I}}}.$
@@
is element of $\mathcal{P}^{1}_{\mathcal{I}}$ and satisfies (58) and (59).
With these preparations, we can prove the Lemmas A.2 and A.1.
Proof A.9 (Proof of Lemma A.2).
For any $l\in\mathbb{N}$, for any ^$\hat{\mathcal{I}}\subset\mathbb{N}_{l}$ and for any $i\in\mathcal{I}$, let =^$\mathcal{I}=\hat{\mathcal{I}}^{-i}=\{s-i|s\in\hat{\mathcal{I}}\text{ with }s>i\}$ and $\{x_{1},\dots,x_{n_{\mathcal{I}}}\}=\{s\in\hat{\mathcal{I}}|s\geq i\}$. Then Lemma A.5 implies the statement for $i\neq 0$ and Lemma A.7 for $i=0$.
Proof A.10 (Proof of Lemma A.1).
The statement follows from Lemma A.2 with the Ansatz
$p({\bm{q}})=\hat{p}\left(\sum_{j=1}^{n}q_{j}\right)$ for all ${\bm{q}}\in\mathbb{R}^{n}$.
A.4 Gappy multivariate polynomial interpolation
@@
Proof A.11 (Proof of Theorem A).
Let $i_{1}<i_{2}<\dots i_{n_{\mathcal{I}}}$ be the elements of $\mathcal{I}$.
Because of Theorem A.2, we can construct polynomials $p_{i_{k}}\in\mathcal{P}^{n}_{\{i_{k}\}}$ sequentially for $k={n_{\mathcal{I}}},{n_{\mathcal{I}}}-1,\dots,2,1$ such that
(72) $p_{i_{k}}({\bm{q}}_{j})=b_{j}-\sum_{s=k+1}^{{n_{\mathcal{I}}}}p_{i_{s}}({\bm{q}}_{j}),\hskip 10.00002ptj=1,2,\dots,{n_{\mathcal{I}}},$
for the interpolation nodes ${\bm{q}}_{j}$ in $\mathcal{X}^{i_{k}}$ and the interpolation values $b_{j}$ prescribed in (39).
Because of Lemma A.1, we can construct polynomials $\tilde{p}_{i_{k}}\in\mathcal{P}^{n}_{\mathcal{I}^{-i_{k}}}$ such that
@@
By multiplying each of these polynomials $p_{i_{k}}\in\mathcal{P}^{n}_{\{i_{k}\}}$ with the corresponding polynomial $\tilde{p}_{i_{k}}\in\mathcal{P}^{n}_{\mathcal{I}^{-i_{k}}}$, we obtain polynomials $\bar{p}_{i_{k}}:=p_{i_{k}}\tilde{p}_{i_{k}}\in\mathcal{P}^{n}_{\mathcal{I}}$ that satisfy
[…]
Finally, we get the desired $p\in\mathcal{P}^{n}_{\mathcal{I}}$ by summing up $p({\bm{x}})=\sum_{k=1}^{n_{\mathcal{I}}}\bar{p}_{i_{k}}({\bm{x}})$. For any fixed $k^{*}\in\{1,\dots,{n_{\mathcal{I}}}\}$, we find for all ${\bm{q}}_{j}\in\mathcal{X}^{i_{k^{*}}}$,
(77) $p({\bm{q}}_{j})$ $=\sum_{k=1}^{n_{\mathcal{I}}}\bar{p}_{i_{k}}({\bm{q}}_{j})=\sum_{k=k^{*}}^{{n_{\mathcal{I}}}}\bar{p}_{i_{k}}({\bm{q}}_{j})$
because of (76), and
(78) $\sum_{k=k^{*}}^{{n_{\mathcal{I}}}}\bar{p}_{i_{k}}({\bm{q}}_{j})=\bar{p}_{i_{k^{*}}}({\bm{q}}_{j})+\left(\sum_{k=k^{*}+1}^{{n_{\mathcal{I}}}}\bar{p}_{i_{k}}({\bm{q}}_{j})\right)=\left(b_{j}-\sum_{k=k^{*}+1}^{{n_{\mathcal{I}}}}\bar{p}_{i_{k}}({\bm{q}}_{j})\right)+\left(\sum_{k=k^{*}+1}^{{n_{\mathcal{I}}}}\bar{p}_{i_{k}}({\bm{q}}_{j})\right)=b_{j},$
because of (75), so $p$ indeed satisfies (39).

### v2

@@ 3 Exact operator inference
In this work, we propose a novel method to generate snapshots that is a priori guaranteed to result in a full rank matrix ${\mathbf{P}}$. Our method has two key ingredients. First, in Section 3.1, we leverage the knowledge about how the matrix ${\mathbf{P}}$ depends on the ROM states $\breve{\bm{x}}_{k}$ and input signals ${\bm{u}}_{k}$ via (12) and (7) to find pairs of ROM states $\bar{\bm{x}}_{s}\in\mathbb{R}^{n}$ and input signals $\bar{\bm{u}}_{s}\in\mathbb{R}^{N_{u}}$, $s=1,\dots,{n_{f}}$ that ensure full rank of ${\mathbf{P}}$. Second, in Section 3.2, we generate time derivative snapshots $\dot{\bar{\bm{x}}}_{s}$ corresponding to these rank-ensuring pairs $(\bar{\bm{x}}_{s},\bar{\bm{u}}_{s})$ by simulating ${n_{f}}$ trajectories for only a single time step each, with initial conditions $\bar{\bm{x}}_{s}$ and input signals $\bar{\bm{u}}_{s}$.
@@ 4.2 Ice sheet model
For ROM dimension $n=7$, the proposed method performs ${n_{f}}=3087$ FOM time steps. This number is larger than the $2001$ FOM time steps with re-projection reported in [2]. However, the authors also report that this budget of snapshots is depleted before the inference of the ROM operators of dimension $n=6$ is completed. Hence, the ROM operators of dimensions $n=6$ and $n=7$ are partially computed with regularized operator inference and snapshot data polluted with non-Markovian terms. Consequently, the inferred operators of dimension $n=6$ and $n=7$ are not reconstructed exactly. The operators of dimension $n\leq 5$ might be reconstructed exactly, but the used $2001$ snapshots are more than the ${n_{f}}=365$ snapshots used by our method for dimension $n=5$.
@@ Appendix A Proof of full rank of matrix ${\mathbf{P}}$
For any $n\in\mathbb{N}^{+}$, $l\in\mathbb{N}_{0}$, for any degree set $\mathcal{I}\subset\mathbb{N}_{l}$, there exists a unique polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}}$ such that
@@
A.1 Outline of proof
The special case of degree sets without gaps, (i.e. $\mathcal{I}=\mathbb{N}_{l}$), is covered by an existing result on existence and uniqueness of an interpolating multivariate polynomial [17] described in Section A.2. In Section A.3, we use this result to prove Theorem A for the special case that $\mathcal{I}$ contains only one element. I.e., we can find for all $i\in\mathcal{I}$ a homogeneous polynomial $p_{i}\in\mathcal{P}^{n}_{\{i\}}$ that interpolates in the nodes in $\mathcal{X}^{i}$. We use this result as base case in a proof by induction in Section A.5 to prove the general case where $\mathcal{I}$ contains multiple elements. There, the induction hypothesis states that Theorem A holds for a fixed degree set ${\mathcal{I}^{-}}$, i.e., there exists a unique polynomial $p^{-}$ that interpolates in the interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$. In the induction step, we show that Theorem A also holds for the new degree set :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ for any ${i^{*}}\in\mathbb{N}_{0}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$. For this purpose, we add a new polynomial $p$ that has two key properties. First, $p$ attains suitable values in the new interpolation nodes in $\mathcal{X}^{{i^{*}}}$ such that the resulting polynomial $p^{-}+p$ attains the prescribed interpolation values $b_{j}$. Second, $p$ attains 0 in the old interpolation nodes in $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$ to not interfere with the interpolation property of $p^{-}$ in these nodes. In Section A.4, we show that such a polynomial $p$ exists in $\mathcal{P}^{n}_{\mathcal{I}}$.
A.2 Multivariate polynomial interpolation without gaps
@@
A.3 Homogeneous multivariate polynomial interpolation
@@
for the prescribed interpolation values $f_{i}$ and the interpolation nodes $\hat{\bm{q}}_{i}$ in $B(l,n-1)$.
@@
A.4 Specific gappy multivariate polynomial interpolation
As explained in Section A.1, the proof of Theorem A in Section A.5 involves a polynomial $p\in\mathcal{P}^{n}_{\mathcal{I}}$ with :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ such that
(54) $p({\bm{q}}_{j})$ $=0,\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i}\text{ for all }i\in{\mathcal{I}^{-}}$
(55) $p({\bm{q}}_{j})$ $=f_{j},\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i^{*}},$
for prescribed interpolation values $f_{j}\in\mathbb{R}$. To show that such a polynomial $p$ exists, we make the Ansatz
(56) $p({\bm{q}})=p_{i^{*}}({\bm{q}})\hat{p}\left(\sum_{k=1}^{n}q_{k}\right)\hskip 10.00002pt\forall{\bm{q}}\in\mathbb{R}^{n},$
with $p_{i^{*}}\in\mathcal{P}^{n}_{i^{*}}$ that satisfies (55) according to Theorem A.3, and a univariate polynomial $\hat{p}$ that solves the specific interpolation problem with interpolation values 0 and 1 described in the following lemma.
@@ A.4 Specific gappy multivariate polynomial interpolation
For any $l\in\mathbb{N}_{0}$, for any :=$\mathcal{I}:=\{{i^{*}}\}\,\cup\,{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, there exists a polynomial $p\in\mathcal{P}^{1}_{{\mathcal{I}^{-{i^{*}}}}}$ with ${\mathcal{I}^{-{i^{*}}}}:=\{i-{i^{*}}\big{|}i\in\mathcal{I}\}$ such that
(57) $p(i)$ $=0,\hskip 10.00002pt\text{ for all }i\in{\mathcal{I}^{-}}\text{ and }$
(58) $p({i^{*}})$ $=1.$
To prove this lemma, we show in Lemma A.2 that for fixed $l\in\mathbb{N}_{0}$, the existence of solutions to such specific interpolation problems with interpolation values 0 and 1 as described above implies the existence of solutions to the general interpolation problems with arbitrary interpolation values. We use this result to prove Lemma A.1 with a case discrimination: In Lemma A.4, we consider the case ${i^{*}}\neq 0$ and in Lemma A.6 the case ${i^{*}}=0$.
@@
for arbitrary interpolation values $f_{j}\in\mathbb{R}$.
Proof A.3.
@@
Proof A.5.
The proof proceeds by induction along $l$.
  • Base case: For $l=0$, $p(x)=1$ satisfies the interpolation properties for the only existing set =$\mathcal{I}=\{0\}$.
  • Induction hypothesis: For fixed $l$, there exists $p\in\mathcal{P}^{1}_{\mathcal{I}}$ such that the interpolation properties (59) and (60) hold for any $\mathcal{I}\in\mathbb{N}_{l}$ and any pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}$.
  • Induction step: We consider a set $\mathcal{I}\subset\mathbb{N}_{l+1}:=\{0,1,\dots,l,l+1\}$ and the pairwise distinct, positive nodes $x_{1},\dots,x_{n_{\mathcal{I}}}\in\mathbb{R}^{+}$. If $l+1\notin\mathcal{I}$, then the induction hypothesis directly applies. Otherwise, if $l+1\in\mathcal{I}$, then the induction hypothesis holds for :=$\mathcal{I}^{-}:=\mathcal{I}\setminus\{l+1\}$ and the nodes $x_{1},\dots,x_{{n_{\mathcal{I}}}-1}$. Hence, Lemma A.2 implies that there exists a polynomial $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
    for arbitrary interpolation values $f_{i}\in\mathbb{R}$. Consequently, there exists $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ such that
@@
    satisfies (59), i.e., $x_{1},\dots,x_{{n_{\mathcal{I}}}-1}$ are roots of $\hat{p}$. By Descartes’ rule of signs, $\hat{p}$ has at most ${n_{\mathcal{I}}}-1$ positive roots, hence $\hat{p}(x_{n_{\mathcal{I}}})\neq 0$. Consequently, we can normalize $\hat{p}$ to find the desired polynomial,
@@
(71) $\bar{p}(x_{j})=-1\hskip 10.00002pt\text{ for }j=1,\dots,{n_{\mathcal{I}}}-1.$
@@
is element of $\mathcal{P}^{1}_{\mathcal{I}}$ and satisfies (59). Because $\bar{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-}}$ with $0\notin\mathcal{I}^{-}$, $x$ is a factor of $\bar{p}$, so $\bar{p}(0)=0$. Consequently $p$ also satisfies (60).
With these preparations, we can prove Lemma A.1.
Proof A.8 (Proof of Lemma A.1).
For any $l\in\mathbb{N}$, for any ^=$\hat{\mathcal{I}}=\{{i^{*}}\}\cup{\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with ${i^{*}}<\min{\mathcal{I}^{-}}$, let =$\mathcal{I}=\{i-{i^{*}}\big{|}i\in\hat{\mathcal{I}}\}$ and $\{x_{1},\dots,x_{n_{\mathcal{I}}}\}=\hat{\mathcal{I}}$. Then Lemma A.4 implies the statement if ${i^{*}}\neq 0$ and Lemma A.6 if ${i^{*}}=0$.
A.5 Gappy multivariate polynomial interpolation
@@
Proof A.9 (Proof of Theorem A).
The proof is performed by induction along the number of elements in $\mathcal{I}$.
  • Base case: Let =$\mathcal{I}=\{i\}$ for any $i\in\mathbb{N}_{l}$ with arbitrary $l\in\mathbb{N}_{0}$, then Theorem A.3 implies the existence and uniqueness of a polynomial $p\in\mathcal{P}^{n}_{\{i\}}$ that satisfies the interpolation property (39) in the interpolation nodes in $\mathcal{X}^{i}$.
  • Induction hypothesis: Let for a fixed degree set ${\mathcal{I}^{-}}\subset\mathbb{N}_{l}$ with arbitrary $l\in\mathbb{N}_{0}$ exist a unique polynomial $p^{-}\in\mathcal{P}^{n}_{\mathcal{I}^{-}}$ that satisfies the interpolation property (39) in the interpolation nodes $\cup_{i\in{\mathcal{I}^{-}}}\mathcal{X}^{i}$.
  • Induction step: We consider the degree set :=$\mathcal{I}:=\{{i^{*}}\}\cup{\mathcal{I}^{-}}$ with arbitrary ${i^{*}}\in\mathbb{N}_{l}$. Without loss of generality, we can assume ${i^{*}}<\min{\mathcal{I}^{-}}$. As described in Section A.4, we can construct a polynomial
    (73) $p({\bm{q}})=p_{i^{*}}({\bm{q}})\hat{p}\left(\sum_{k=1}^{n}q_{k}\right)\hskip 10.00002pt\forall{\bm{q}}\in\mathbb{R}^{n},$
    with $p_{i^{*}}\in\mathcal{P}^{n}_{i^{*}}$ and $\hat{p}\in\mathcal{P}^{1}_{\mathcal{I}^{-{i^{*}}}}$, so $p\in\mathcal{P}^{n}_{\mathcal{I}}$. Because of Lemma A.1 and $\sum_{k=1}^{n}q_{k}=i$ for all ${\bm{q}}\in\mathcal{X}^{i}$ for any $i\in\mathbb{N}_{0}$ , we can choose $\hat{p}$ such that
    (74) $\hat{p}\left(\sum_{k=1}^{n}q_{k}\right)$ $=0,\hskip 10.00002pt\text{ for all }{\bm{q}}\in\mathcal{X}^{i}\text{ for all }i\in{\mathcal{I}^{-}}$
    (75) $\hat{p}\left(\sum_{k=1}^{n}q_{k}\right)$ $=1,\hskip 10.00002pt\text{ for all }{\bm{q}}\in\mathcal{X}^{i^{*}},$
    and $p_{i^{*}}$ such that
@@
    with the interpolation values $b_{j}$ prescribed in (39). Consequently, the polynomial $p$ satisfies
    (77) $p({\bm{q}}_{j})$ $=0,\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i}\text{ for all }i\in{\mathcal{I}^{-}}\;\text{ and }$
    (78) $p({\bm{q}}_{j})$ $=b_{j}-p^{-}({\bm{q}}_{j}),\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i^{*}}.$
    As a result, the polynomial $p^{-}+p$ satisfies
    (79) $p^{-}({\bm{q}}_{j})+p({\bm{q}}_{j})$ $=b_{j}+0=b_{j},\hskip 10.00002pt\text{ for all }{\bm{q}}_{j}\in\mathcal{X}^{i}\text{ for all }i\in{\mathcal{I}^{-}}\;\text{ and }$
@@
    and hence satisfies the interpolation property (39). Since the corresponding interpolation matrix is square, existence of a solution also implies uniqueness.